40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1498     Entry time: Mon Apr 20 05:18:42 2009
Author: Yoichi 
Type: Configuration 
Category: Locking 
Subject: FM6 and FM10 of LSC-MC were restored 
During tonight's locking work, I realized that FM6 and FM10 (both resonant gains around 20Hz) were actually activated by cm_step.
So I restored those filters from the svn history.
Instead, I removed a bunch of unused filters from LSC-DEMOD and LSC-DEMOD_A (moving zero filters) to off load c1lsc.

As for the locking itself, the DARM loop becomes unstable at around arm power = 30. I may have to add a filter to give a broader phase bubble.
ELOG V3.1.3-